# Spatial Knowledge Graph-Guided Multimodal Synthesis

### v1

@@ 4.2 Main Results
The fine-tuned LLaVA-1.6 model achieves an average score of 50.0 (+1.8), while the fine-tuned Llama-3.2-vision model reaches 51.7 (+4.0). These improvements confirm that SKG2Data effectively enhances spatial reasoning while maintaining performance in general visual understanding tasks.
4.3 Ablation and Analysis
@@
Figure 6: Case analysis: On spatially related datasets, the trained model enhances spatial understanding abilities.
Finding 1: Directional knowledge plays a crucial role in enhancing the MLLM’s spatial perception capabilities.
@@
Finding 2: Data synthesized from two types of spatial knowledge, including direction and distance, exhibit generalization ability.
@@
Finding 3: Increasing the number of objects improves the performance of MLLMs in visual detail tasks
@@
Discussion on Knowledge-to-Data Synthesis.
It can be seen that with more data, the performance of the trained model can be further improved. At the same time, it can also be found that high-quality data is important, and a small amount of high-quality training data can significantly improve the performance of models in specific fields, which is in line with the opinion of “less is more” Zhou et al. (2023); Ye et al. (2025). As shown in Figure 6, the trained model exhibits improved accuracy on spatially related questions. The core idea of our work is to synthesize data through automated modeling knowledge to ensure the rationality of the data itself. The data generated through spatial knowledge is consistent with human spatial cognition, ensuring the relevance and reliability of the data. In addition, to synthesize additional data, we can either randomly sample more data from the SKG or expand the scenes to further enrich the SKG. Compared with the method of synthesizing multimodal data using code, simulators, etc., it is more efficient and has better generalization ability.
5 Related Work
Multi-Modal Large Language Model.
The remarkable success of large language models (LLMs) Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023) has driven the development of MLLMs Ye et al. (2023); Liu et al. (2024a); Chen et al. (2023); Meta AI (2024); Gemini Team (2024); Bai et al. (2023); OpenAI (2023, 2024). By incorporating visual backbones, these MLLMs demonstrate strong performance across a wide range of visual-language tasks, significantly improving the processing and understanding of both text and visual information. At the same time, a variety of benchmarks Yu et al. (2023); Guan et al. (2024); Chen et al. (2024) have been developed to evaluate the capabilities of large multimodal models. Additionally, with increasing attention on spatial understanding in MLLMs, specialized benchmarks Kamath et al. (2023); Du et al. (2024); Tong et al. (2024) have been introduced. These benchmarks are specifically designed to tackle the challenges of evaluating spatial comprehension in MLLMs.
Synthetic Data Generation.
Synthetic data has garnered significant attention as a novel solution to the challenges of obtaining large, diverse, and high-quality datasets Chang et al. (2024); Liu et al. (2024b); Long et al. (2024); Kim et al. (2024). Many existing studies primarily rely on pre-existing images Wu et al. (2024), occasionally incorporating additional image information such as captions, bounding boxes, or OCR input Liu et al. (2023a); Zhao et al. (2023a); Zhang et al. (2023). However, another area of research transcends generating textual instruction data and focuses on creating high-quality images through various synthesis techniques rather than relying solely on existing image datasets. For example, REACHQA He et al. (2024) and Multimodal Self-Struct Zhang et al. (2024) use code to precisely synthesize chart images. Additionally, EmbSpatial utilizes simulators to generate images, while SynCLR Tian et al. (2024) and VisMin Awal et al. (2024) employ diffusion models to produce new images. Furthermore, due to the tendency of LLMs to exhibit biases and follow a long-tailed distribution due to their lack of specialization in specific domains Navigli et al. (2023); Fei et al. (2023), there is a growing body of research Feng et al. (2023); Xu et al. (2024); Wang et al. (2024); Yang et al. (2025); Ovadia et al. (2025) that employs knowledge enhancement techniques, such as KG, to generate higher-quality data.
6 Conclusion
@@
In this work, we propose SKG2Data, a novel multimodal synthetic data generation method that leverages Spatial Knowledge Graph to synthesize images and question-answer pairs, addressing the lack of high-quality data for spatial understanding tasks. Using SKG2Data, we create a multimodal instruction dataset and a benchmark, SKG2Data-Holdout, to enhance and evaluate the spatial understanding capabilities of MLLMs. Extensive experiments demonstrate the effectiveness of SKG2Data and analyze the impact of spatial relationship types, dataset scales, and object densities.
Limitations
Noise in synthetic data.
Despite the implementation of automatic filtering, the synthesized data may still contain noise. This noise arises from potential mismatches between images and text in diffusion models, as well as the inherent possibility of hallucinations in LLMs.
Difficult to synthesize data with a large number of objects.
Given the limitations of current diffusion models, generated images can sometimes show problems like distortion and inconsistencies in object type, attributes, quantity, or spatial arrangement. The challenge of creating realistic images grows substantially as the number of objects in the scene increases. In some real-world situations where scenes are packed with many objects, our method may have difficulty accurately depicting these complex scenes.
Appendix A Appendix
Hyperparameter Setting
Fine-tuning method LoRA
Gobal Batch Size 64
[…]
Learning Rate (LR) $51e-$ 5
LoRA $r$ 16
LoRA $\alpha$ 32
LoRA Dropout 0.05
Epoch 2
Mixed precision Enabled
Table 2: Detailed experimental settings.
Figure 7: Training data cases.
Prompt for "Scenes Generation":
You are an AI assistant specializing in imagining the scenes depicted in photos. Your task is to envision a variety of scenes that may appear in different photos. Your response needs to meet the following requirements: 1. Ensure that every scene is safe, accessible to most people, and common in real and daily life scenes. 2. The scenes in the photos should be diverse and non repetitive. Each generated scene typically depicts a facility (usually requiring interaction with multiple objects such as a table or furniture) or a room, or a place or area. Please directly output the generated scenes in the following list format: <In context learning> Now, please generate 160 different scenes. <image>
Prompt for "Objects Generation":
<Wikipedia documents> You are an AI assistant specializing in identifying objects likely to appear in a given scene. Your task is to list as many objects as possible that may be found in that scene. Your response needs to meet the following requirements: 1. Ensure that each object is based on common situations in reality and daily life. Ensure that objects are relatively simple and easy to draw, and once you see their images, you can determine what objects they are. Ensure that the object is safe and harmless. 2. The objects provided should not be duplicated or too similar. Try to describe an object with one word. If there are words with multiple meanings, it is necessary to ensure that they are distinguishable, such as "animal mouse" and "computer mouse". 3. Each scene needs to list as many objects as possible. If there may be people on site, please add ’person’. Here I will give you some examples for reference and use the following template in the format of a list for the response: <In context learning> <image>
Prompt for "Spatial KG Generation (Nodes)":
You are an AI assistant specializing in describing objects in a given scene. Your task is to first select a subset of candidate objects, and then describe the attributes of the selected objects. Your response should meet the following requirements: 1. Select 2 to 4 suitable objects from the given objects to form a possible combination. Note that for some candidate objects that are suitable for repeated appearance, they can be selected repeatedly. Please prioritize selecting more common objects. Do not choose objects that are difficult to represent or recognize in the image. 2. Provide further attribute descriptions for the selected objects, including color, orientation, etc., and describe countable objects as singular. Be careful not to describe objects that do not exist in the real world or are unsuitable for the current scene. The description of each object should be concise, without being overly complex and abstract, and avoid including details that cannot be visualized in the image. For some objects that are difficult to add descriptions, additional descriptions can be omitted. 3. The selected combination of objects can be repeated, but the described attributes need to be different. 4. If there are words with multiple meanings, it is necessary to ensure that they are distinguishable and you need to determine which one is more likely based on the scene and other objects, such as ’animal mouse’ and ’computer mouse’. 5. You need to provide 25 to 50 possible combinations based on the number of objects given. Here I will give you some demonstrations for reference and use the following list template for the response: <In context learning> <image>
Prompt for "Spatial KG Generation (Triplets)":
You are an AI assistant specializing in describing the positional relationship between objects in a scene. Your task is to describe the positional relationship of the given objects. Your response needs to meet the following requirements: 1. Be careful not to describe positional relationships that do not exist in the real world. When describing the positional relationship between two objects, do not involve other objects. At most one positional relationship triplet is generated between every two objects. 2. There may be two types of positional relationships between two objects: directional relationships and distance relationships. Directional relationships include above, below, left, right, in front of, behind, etc. Distance relationships include close, far, near, etc. Separate two positional relationships with commas when generating them simultaneously. For situations with a small number of objects, prioritize generating directional relationships. Here I will give you some demonstrations for reference and use the following template in the format of a list for the response: <In context learning> <image>
Table 3: The prompt examples that we use in the first module.
Prompt for "Image Data Generation":
You are an intelligent bounding box generator. I will provide you with the scene, objects and their spatial relationships in a photo. Your task is to generate the bounding boxes for the given objects, along with a background prompt describing the scene and a prompt describing the photo. The images are of size 512x512. The top-left corner has coordinate [0, 0]. The bottom-right corner has coordinnate [512, 512]. The bounding boxes should not overlap or go beyond the image boundaries. Each bounding box should be in the format of (object name, [top-left x coordinate, top-left y coordinate, box width, box height]) and should not include more than one object. Do not put objects that are already provided in the bounding boxes into the background prompt. Do not include non-existing or excluded objects in the background prompt. Use "A realistic scene" as the background prompt if no background is given in the prompt. There may be multiple spatial relationships between objects, and special attention should be paid to using bounding boxes to represent these spatial relationshipss. Please refer to the example below for the desired format. <In context learning> <image>
Prompt for "Image Data Filtering and Textual Data Generation":
You are an AI assistant specializing in evaluating the quality of a given image and generating question and answer pairs related to the image. Your task is to first determine whether the given image should be saved based on its quality and whether it matches the descriptions, and then continue generating question and answer pairs if it is determined to be saved. Descriptions of the image will provide information about the objects present in the image, their quantity, their attributes, their positional relationships, their bounding boxes (object name, [top-left x coordinate, top-left y coordinate, box width, box height]), and objects that are absent from the image. You need to follow the steps below to complete the task: Step 1. You need to ensure the clarity, quality, and authenticity of the image. You need to verify if there are any unrealistic or unreasonable parts in this image, such as clearly unrealistic looking humans or animals, and if there are any indistinguishable objects. Step 2. You need to ensure that the image contains the specified objects and does not contain any excluded objects. Specified objects are given in the form of dict, and excluded objects are presented in a list format. Please note that the image should accurately reflect the attributes of the specified object, such as Orientation and Direction, State and Condition, Color and Appearance, Structural and Physical Characteristics, Viewpoint and Perspective. Step 3. You need to ensure that the objects in the image have a positional relationship with the matching descriptions. The positional relationships are given in the form of a list, and the objects correspond to the dict. Step 4. If the image matches step 1 and basically matches steps 2 and 3, please answer ’Yes’ and proceed to the next steps to generate question answer pairs. Otherwise, please answer ’No’ and do not proceed to the next steps. Step 5. You need to generate as many question answer pairs as possible based on the correct parts of these descriptions and information that can be confidently determined in the given image, while ensuring quality and diversity. For Step 5, the following requirements must be met: 1. Please note not to mention any information related to the image descriptions in your question-answer pairs, as the image description will not be provided in future tasks using these question-answer pairs. 4. Use the template in the format of a JSON list to generate question and answer pairs without any other information. Below are reference responses for saving and not saving this image: <In context learning> <image>
Prompt for "Textual Data Filtering":
You are an AI assistant specializing in reviewing question-answer pairs about a given image. Your task is to find the correct question-answer pairs based on the given image. Your response needs to meet the following requirements: 1. You need to determine whether the answers to the questions are correct based on the given images, and save the correct question-answer pairs. 2. Use a JSON list format template with the same format as the input to generate correct question-answer pairs without generating any additional information. Here is the question and answer about the given image: <In context learning> <image>
Table 4: The prompt examples that we use in the second module.

### v2

@@ V-B Main Results
Many current works have found that training on datasets with specific domains or abilities can affect the performance of models in other domains and general domains, leading to an overall performance declines [64, 65, 66]. For our method, the fine-tuned LLaVA-1.6 model achieves an average score of 50.0 (+1.8), while the fine-tuned Llama-3.2-vision model reaches 51.7 (+4.0). These improvements confirm that SKG2Data effectively enhances spatial reasoning while maintaining performance in general visual understanding tasks.
Figure 7: Cases analysis: In spatially related datasets, the trained model enhances spatial understanding abilities.
V-C Ablation and Analysis
@@
Finding 1. Directional knowledge plays a crucial role in enhancing the MLLM’s spatial perception capabilities
@@
Finding 2. Data synthesized from two types of spatial knowledge, including direction and distance, exhibit generalization ability
@@
Finding 3. Increasing the number of objects improves the performance of MLLMs in visual detail tasks
@@
Noise and Error Analysis
Despite the implementation of automatic filtering, the synthesized data may still contain noise. This noise arises from potential mismatches between images and text in diffusion models, as well as the inherent possibility of hallucinations in MLLMs. In addition, due to the inherent limitations of current diffusion models, generated images often show problems like distortion and inconsistencies in object type, attributes, quantity, or spatial arrangement. Creating realistic images becomes much harder as the SKG becomes more complex. Specifically, when the SKG grows larger—with more nodes representing objects and more edges encoding their relationships—the image generation task becomes significantly more difficult. Therefore, the various SKG graph structures currently sampled for synthesizing training data are not too complex. In some real-world situations where scenes are packed with many objects and complex spatial relationships, our method may have difficulty accurately generating these complex data.
Figure 8: The mechanism of changes in the LLaVA model before and after training. We use two cases to calculate the probability of tokens representing various options at each layer of the transformer, which reveals that the trained model improves the probability of correct options in the last few layers, especially in the last layer.
Discussion on Knowledge-to-Data Synthesis
As shown in our experiments, increasing the volume of training data generally leads to improved model performance. Moreover, we observe that high-quality data is important and a small amount of high-quality training data can significantly improve the performance of models in specific fields, which is in line with the opinion of “less is more” [67, 68]. As illustrated in the three cases of Figure 7, the trained model demonstrates significantly improved accuracy in spatially related datasets. To further investigate how the LLaVA model evolves during training, we select two types of case: one that the model answers correctly even before training and another that it answers correctly only after training. We analyze internal representations [69, 70, 71] in different layers by processing these inputs through the model. An LM-head attached to each layer produces logits for the first generated token, which are then converted into probabilities via softmax. As shown in Figure 8, the token probability distribution remains largely unchanged in the previous transformer layers. However, in the last few transformer layers, especially in the last layer, the probability of correct token is rapidly improved.
The core idea of our work is to synthesize data through automated modeling knowledge to ensure the rationality of the data itself. The data generated through spatial knowledge is consistent with human spatial cognition, ensuring the relevance and reliability of the data. In addition, to synthesize additional data, we can either randomly sample more data from the SKG or expand the scenes to further enrich the SKG. Most existing simulators [72, 73, 74, 75, 76, 77, 78] are designed for specific environments, and image synthesis via code is largely confined to mathematical visualizations or web-based content. Consequently, such approaches are unable to effectively generate multimodal spatial data beyond these limited contexts. In contrast, our method supports effective data generation across a wide variety of scenarios.
VI Conclusion
@@
In this work, we propose SKG2Data, a novel multimodal synthetic data generation method that leverages Spatial Knowledge Graph to synthesize images and question-answer pairs, addressing the lack of high-quality data for spatial understanding tasks. Using SKG2Data, we create a multimodal instruction dataset and a benchmark, SKG2Data-Holdout, to enhance and evaluate the spatial understanding capabilities of MLLMs. Extensive experiments analyze the effectiveness of SKG2Data and the impact of spatial relationship types, dataset scales, and object densities. Our results show that SKG2Data significantly improves MLLMs’ performance on spatial reasoning tasks. Furthermore, the automatic construction of Spatial Knowledge Graphs and data synthesis pipeline offers a flexible framework that can be extended to various domains requiring precise spatial representations.
Acknowledgments
We would like to express gratitude to the anonymous reviewers for their hard work and kind comments. This work was supported by the National Natural Science Foundation of China (No. NSFCU23B2055, No. 62576307, No.62506128), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China under Grant (No. LRG25F030003 and LQN25F020023), Yongjiang Talent Introduction Programme (2021A-156-G), Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG.